A square ABCD and PQ are the midpoints of BC and CD. If AP=a and AQ=b .find in terms of a and b the directed line segment AB


1
Expert's answer
2022-03-04T05:41:10-0500

The square is presented on the picture:



Denote by "x" the length of the side of the square. "a=|AP|=\sqrt{|AB|^2+|BP|^2}=\sqrt{x^2+\frac{1}{4}x^2}=x\frac{\sqrt{5}}{2}". "b=|AQ|=x\frac{\sqrt{5}}{2}" . Thus, we received that "a=b=x\frac{\sqrt{5}}{2}". Thus, we get: "x=\frac{2a\sqrt{5}}{{5}}=\frac{2b\sqrt{5}}{{5}}". The length of "AB" is "\frac{2a\sqrt{5}}{{5}}=\frac{2b\sqrt{5}}{{5}}"

Answer: "\frac{2a\sqrt{5}}{{5}}=\frac{2b\sqrt{5}}{{5}}"